# Reducing Energy Bloat in Large Model Training

### v2

@@ 3.2. Perseus Architecture
Figure 4 shows the high-level architecture of Perseus. Perseus is split into a framework- and hardware-agnostic server and a framework-integrated and device-specific client. The server pre-characterizes the “iteration time–energy” Pareto frontier (§4) and caches them for fast lookup. The client profiles pipeline computations online during training and realizes the energy schedule by changing the GPU’s frequency during runtime (§5).
@@
Training Lifecycle Using Perseus.
In Perseus, a training job is primarily specified by its computation DAG for one training iteration. When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy of each forward and backward computation on each frequency while training is running.
Upon receiving profile results, 2 the server characterizes the “iteration time–energy” Pareto frontier (§4), and Pareto-optimal energy schedules are saved in a lookup table indexed by $T^{\prime}$. Then, 3 the energy schedule corresponding to the shortest iteration time is deployed to the client. Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework (§5).
During training, 4 the training infrastructure (e.g., datacenter rack power/temperature manager) notifies the Perseus server of a straggler and its degree of slowdown. The server then 5 quickly reacts to this by looking up the Pareto-optimal energy schedule corresponding to the anticipated straggler iteration time, and deploys it to the client.
4. Characterizing the Pareto Frontier
In this section, we describe our algorithm to efficiently obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail. We first formulate the problem, show that it is NP-hard, and provide a relaxation (§4.1). Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3). Finally, we extend our algorithm to support 3D/hybrid parallelism, constant-time operations, and diverse pipeline schedules (§4.4).
@@ 4.1. Problem Formulation
where $P_{\textrm{blocking}}$ is the power consumption of the GPU when it is blocking on communication, $N$ is the number of pipeline stages, and $t_{i}(f_{i})$ and $e_{i}(f_{i})$ are the time and energy consumption of computation $i$ with frequency $f_{i}$, respectively.
As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the “iteration time–energy” Pareto frontier ($\textrm{Time}(F)$ vs. 1+ 2) at $T_{\textrm{opt}}$ and find $F_{\textrm{opt}}$ where the two lines intersect. Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline only depends on the straggler’s iteration time $T^{\prime}$ in 2, which is only a constant upwards shift of the frontier. Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs. 1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$. Thus, we define
@@
and characterize the Pareto frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
Finding the Pareto Frontier.
Finding one point on the Pareto frontier with iteration time $T$ is equivalent to solving the following optimization problem:
@@
The complete Pareto frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is because the possible choices of GPU frequencies are discrete. However, if choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ). This is akin to integer linear programs becoming tractable when relaxed to linear programs.
The transform from the original problem to the relaxed version is done by fitting a continuous exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements for each forward and backward computation. We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realizes most of the opportunity for savings. Solving the relaxed problem returns the time and energy consumption planned for each computation in the pipeline, or the energy schedule. Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$ (Section 4.3).
4.2. Iteratively Discovering the Pareto Frontier
Figure 5. Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
Although our relaxed problem is no longer NP-hard, solving it for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient. Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate its neighbor energy schedule on the Pareto frontier? Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, discovering every Pareto-optimal energy schedule.
@@ 4.2. Iteratively Discovering the Pareto Frontier
We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) would be incorrect. That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
@@
Output: Set of Pareto-optimal energy schedules $\mathcal{P}$
@@
7 $p\leftarrow$ Minimum energy for all computations
8 $\mathcal{P}\leftarrow\{p\}$
@@
10while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
@@
11       $p\leftarrow$ GetNextPareto($\mathcal{G}$, $p$, $\tau$)
12       $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
@@
$\triangleright$ Every schedule in $\mathcal{P}$ is Pareto-optimal
16 return $\mathcal{P}$
@@
Algorithm 1 Iteratively Discovering the Pareto Frontier.
Algorithm 1 provides an overview of our optimization process. First, the energy schedule with the minimum energy consumption is constructed by planning every computation to run with minimum energy (line 1). Starting from there, the iteration time of the schedule is iteratively reduced by unit time $\tau$ while incurring minimal energy increase (line 1; Section 4.3). This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process is Pareto-optimal.
4.3. Finding the Neighbor Pareto-Optimal Schedule
Figure 6. A simplified example of how to reduce iteration time by unit time $\tau$. Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2). Then the Critical DAG ( 3) is obtained by considering every and only the computations on the critical path. Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$. Cut A and Cut B are two examples of valid s-t cut ( 4). Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
In this section, we describe our core subroutine GetNextPareto (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
@@ 4.3. Finding the Neighbor Pareto-Optimal Schedule
We have shown that an s-t cut represents a way to speed up the entire DAG by $\tau$. But speeding up computations increases energy. Then, a natural question is, which cut brings the smallest energy increase?
We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity. Then, our problem reduces to minimum cut, which we can solve with maximum flow. After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining a new Pareto-optimal energy schedule. Appendix D provides details with mathematical expressions for flow capacities.
@@
Table 2. The minimal set of Perseus client and server APIs that require implementation. One client process manages each accelerator. The type parameter should be either "forward" or "backward". On GPUs, the “speed” control knob is the SM frequency. The server API will issue an HTTP/RPC request to the server. The id parameter is the ID of an accelerator.
@@
Our optimization algorithm has polynomial runtime. Let $N$ and $M$ denote the number of stages and microbatches, respectively. Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$. While for general DAGs the total number of steps is known to be exponential to the size of the DAG (skutella-thesis, ), we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$. See Appendix E for proof.
In reality, commonly used number of stages ($N$) is 4 to 8 (at most tens) to reduce pipeline bubble ratio (megatronlm-sc21, ; mlperf_training_31, ). Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far less have been reported even for high-performance settings (mlperf_training_31, ). As such, algorithm runtime is practically negligible (§6.5), especially given that large model training easily takes weeks or months (patterson2021carbon, ).
@@
Current Pareto-optimal energy schedule $p$
@@
Output: Pareto-optimal schedule w/ reduced time $p^{\prime}$
@@ 12             Remove $i$ from $\mathcal{G}$
18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$)
@@ 20 Modify duration of $\forall i$ in $S-T$ cut by $\tau$
22 $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
@@
Algorithm 2 GetNextPareto: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
@@ 4.4. Generalizations
Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latency. For instance, loading and copying input data into VRAM or communication over slower links can take considerable latency. However, the time and energy consumption of these operations are not affected by the GPU’s frequency. Perseus can take constant-time operations into account during planning by viewing them as a node with only one frequency choice.
@@ 5. Implementation
The Perseus server and client are implemented in Python. Perseus can optimize the energy consumption for any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented.
As a reference, we have integrated the Perseus client with Merak (merak-tpds23, ), which marries high-performance tensor parallelism of Megatron-LM (megatronlm-github, ) and the generic pipeline execution engine of DeepSpeed (deepspeed-github, ). We provide an example of what is looks like to integrate the client with a training engine in Appendix F. Activation recomputation (checkpointing-arxiv16, ) is enabled to allow large batch sizes to fit in GPUs.
Online Time/Energy Profiler.
The client profiles the time and energy of each forward and backward computation in vivo at the beginning of the training job to take measurements representative of real training. While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs for measurement. The GPU’s frequency is profiled from the highest to the lowest at iteration granularity and stopped when energy consumption increases. Beyond that point, frequencies will consume more time and energy, making them suboptimal.
@@ 6. Evaluation
  • Energy bloat reduction was possible because Perseus can enumerate efficient energy schedules on the “iteration time–energy” Pareto frontier (§6.4).
@@ 6.1. Experimental Setup
We report energy reduction and slowdown of a training iteration (%), relative to using all maximum GPU frequencies. In most cases slowdown is close to zero, in which case energy and power reductions coincide. Reducing only extrinsic bloat is not possible, because Perseus reduces extrinsic bloat while keeping intrinsic bloat low as it slows down non-straggler pipelines. Therefore, we report (1) intrinsic bloat reduction without stragglers and (2) intrinsic + extrinsic bloat reduction with stragglers.
@@
  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).

### v3

@@ 3.2. Perseus Architecture
Perseus’s architecture is shown in Figure 4. Perseus consists of a framework- and accelerator-agnostic server and a framework-integrated and accelerator-specific client. The server is a cluster-wide singleton. For various training jobs, the server pre-characterizes the time–energy frontier of one iteration (§4) and caches energy schedules for fast lookup. The client profiles pipeline computations online during training and realizes energy schedules by setting the GPU’s frequency during runtime (§5).
@@
Training Lifecycle.
For the Perseus server, a training job is primarily specified by its computation DAG for one training iteration. When the job begins execution, 1 the Perseus client invokes its online time–energy profiler (§5) to measure the time and energy of each forward and backward computation on each supported frequency. Profiling is done in vivo during the initial tens of training iterations.
Upon receiving profiling results, 2 the server begins asynchronously characterizing the time–energy frontier (§4) while training continues. When characterization finishes, energy schedules on the frontier are saved in a lookup table indexed by $T^{\prime}$. Then, 3 the energy schedule corresponding to $T_{\min}$ is deployed to the client. Energy schedules are realized by the client’s asynchronous frequency controller, integrated into the training framework (§5).
During training, 4 the training infrastructure (e.g., datacenter rack power/temperature manager) notifies the Perseus server of a straggler and its degree of slowdown. The server then 5 quickly reacts to this by looking up the energy schedule corresponding to the anticipated straggler iteration time (the one with iteration time $T_{\textrm{opt}}$), and deploys it to the client.
4. Characterizing the Time–Energy Frontier
In this section, we describe our algorithm to efficiently obtain the time–energy tradeoff frontier for a training pipeline in detail. We first formulate the problem, show that it is NP-hard, and describe a relaxed version (§4.1). Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3). Finally, we extend our algorithm to support 3D/hybrid parallelism, constant-time operations, and diverse pipeline schedules (§4.4).
@@ 4.1. Problem Formulation
where $P_{\textrm{blocking}}$ is the power consumption of the GPU when it is blocking on communication, $N$ is the number of pipeline stages, and $t_{i}(f_{i})$ and $e_{i}(f_{i})$ are the time and energy consumption of computation $i$ with frequency $f_{i}$, respectively. [^5] An assumption here is that $P_{\textrm{blocking}}$ is constant, as a GPU blocking on communication is busy-looping inside a NCCL kernel without heavy computation utilization.
As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the time–energy tradeoff frontier ($\textrm{Time}(F)$ vs. 1+ 2) at $T_{\textrm{opt}}$ and find $F_{\textrm{opt}}$ where the two lines intersect. Equation 3 shows that the time–energy frontier of a pipeline depends on the straggler’s iteration time $T^{\prime}$ only in the second term 2, which is merely an upward shift of the frontier. Therefore, if we characterize the tradeoff frontier of $\textrm{Time}(F)$ vs. 1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any straggler iteration time $T^{\prime}$. Thus, we define
@@
and characterize the frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
Finding the Time–Energy Frontier.
Finding one point on the Pareto-optimal tradeoff frontier with iteration time $T$ is equivalent to solving the following optimization problem:
@@
The complete Pareto-optimal tradeoff frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem that will yield a nearly Pareto-optimal frontier.
One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is in turn because GPUs only support discrete frequencies (e.g., in 15 MHz steps and nothing in the middle). However, if frequency choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ). This is akin to integer linear programs becoming tractable when relaxed to linear programs.
The transform from the original problem to the relaxed version is done by fitting a continuous exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements for each forward and backward computation. We choose the exponential function due to its inherent flexibility and natural fit to data (more details in Appendix D). We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realize most of the opportunity for savings. Solving the relaxed problem returns the time and energy consumption planned for each computation in the pipeline, or the energy schedule. Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$.
4.2. Iteratively Discovering the Frontier
Figure 5. Starting from the energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up and iteratively discover the tradeoff frontier.
Now, we first describe our iterative strategy of characterizing the frontier, and dive deeper into one iteration in Section 4.3.
Although our relaxed problem is no longer NP-hard, solving it for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient. Instead, what if we can tweak an existing schedule already on the frontier to generate its neighbor energy schedule on the frontier? Then, we can start from one end of the frontier and trace along to the other end, discovering fine-grained optimized energy schedules.
@@ 4.2. Iteratively Discovering the Frontier
We note that tracing down from the energy schedule that consumes the maximum energy (i.e., Figure 5 black dot) would be incorrect. That schedule is far from optimized because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
@@
Output: Set of optimized schedules $\mathcal{S}$
@@
7 $s\leftarrow$ Minimum energy for all computations
8 $\mathcal{S}\leftarrow\{s\}$
@@
10while $\mathrm{IterationTime}(\mathcal{G},s)>T_{\min}$ do
@@
11       $s\leftarrow$ GetNextSchedule($\mathcal{G}$, $s$, $\tau$)
12       $\mathcal{S}\leftarrow\mathcal{S}\cup\{s\}$
@@
16return $\mathcal{S}$
@@
Algorithm 1 Iteratively discovering the frontier.
Algorithm 1 provides an overview of our optimization process. First, the energy schedule with the minimum energy consumption is constructed by planning every computation to run with minimum energy (line 1). Starting from there, the iteration time of the schedule is iteratively reduced by unit time $\tau$ while incurring minimal energy increase (line 1; Section 4.3). This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process forms our frontier.
4.3. Finding the Neighbor Energy Schedule
Figure 6. A simplified example of how to reduce iteration time by unit time $\tau$. Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2). Then the Critical DAG ( 3) is obtained by considering only the computations on the critical path. Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$. Cut A and Cut B are two examples of valid s-t cuts ( 4). Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
In this section, we describe our core subroutine GetNextSchedule (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
@@ 4.3. Finding the Neighbor Energy Schedule
We have seen that any s-t cut represents a way to speed up the entire DAG by $\tau$. But speeding up computations increases energy. Then, a natural question is, which cut brings the smallest energy increase?
We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges. That is, by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity, we can reduce our problem to minimum cut, which we can solve with maximum flow. After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining the neighbor energy schedule. Appendix E provides details with mathematical expressions for flow capacities.
@@
Table 2. The minimal set of Perseus client and server APIs that require implementation. One client process manages each accelerator. The type parameter should be either "forward" or "backward". On GPUs, the “speed” control knob is the SM frequency. set_straggler is invoked by the infrastructure with the id of an accelerator to notify the server via HTTP/RPC.
@@
Our optimization algorithm has polynomial runtime. Let $N$ and $M$ denote the number of stages and microbatches, respectively. Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$. While for general DAGs the total number of steps is known to be exponential to the size of the DAG (skutella-thesis, ), we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$. See Appendix F for proof.
In reality, commonly used number of stages ($N$) is 4 to 8 (at most tens) to reduce pipeline bubble ratio (megatronlm-sc21, ; mlperf_training_31, ). Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far fewer have been reported even for high-performance settings (mlperf_training_31, ). As such, algorithm runtime is practically negligible (§6.5), especially given that large model training easily takes weeks or months (patterson2021carbon, ).
@@
Current energy schedule $s$
@@
Output: Neighbor schedule with reduced time $s^{\prime}$
@@ 12             Remove $i$ from $\mathcal{G}$
18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $s$)
@@ 20 Modify duration of $\forall i$ in $S-T$ cut by $\tau$
22 $s^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
@@
Algorithm 2 GetNextSchedule: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
@@ 4.4. Generalizations
There are operations in the training pipeline that may take non-trivial latency, other than computation and blocking on communication. For instance, loading and copying input data into VRAM or communication over slower links can take considerable latency. However, the time and energy consumption of these operations are not affected by the GPU’s frequency. Perseus can take constant-time operations into account during planning by viewing them as a node with only one frequency choice.
@@ 5. Implementation
The Perseus server and client are implemented in Python. Perseus can optimize any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented, and the accelerator supports multiple execution speeds that trade off computation time and energy.
As a reference, we have integrated the Perseus client with Merak (merak-tpds23, ), which marries high-performance tensor parallelism of Megatron-LM (megatronlm-github, ) and the generic pipeline execution engine of DeepSpeed (deepspeed-github, ). While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs. We provide an example of what is looks like to integrate the client with a training engine in Appendix G. Activation recomputation (checkpointing-arxiv16, ) is enabled to allow large batch sizes to fit in GPUs.
Profiler.
Accurate profiling is important to our optimization algorithm; inaccurate latency profiles (especially underestimations) may slow down the end-to-end latency of the DAG, whereas inaccurate energy profiles can lead the algorithm to incorrectly select computations to speed up.
Fortunately, the latency of a fixed set of GPU computations, especially with the GPU’s frequency locked, is known to be very stable (nvml, ; clockwork-osdi20, ). Furthermore, to ensure that profiling results are representative of real training, the Perseus client profiles the time and energy of each forward and backward computation at the beginning of the training job in vivo. Each supported GPU frequency is profiled one by one from the highest to the lowest for about five iterations (more if one iteration has less microbatches). After a certain frequency, lower frequencies result in both more time and energy consumed, making them strictly suboptimal compared to higher frequencies. Profiling is terminated at that point.
Finally, we profile $P_{\textrm{blocking}}$ using two GPUs. One GPU blocks on P2P communication and the other sleeps, and we measure the power consumption of the blocking GPU. It is sufficient to profile $P_{\textrm{blocking}}$ once per GPU model.
@@ 6. Evaluation
  • Energy bloat reduction is possible because Perseus can enumerate efficient energy schedules on the time–energy frontier (§6.4).
@@ 6.1. Experimental Setup
We report GPU energy reduction and slowdown of a training iteration (%) relative to using all maximum GPU frequencies. In most cases slowdown is close to zero, in which case energy and average power reductions coincide. Reducing only extrinsic bloat is not possible, because Perseus reduces extrinsic bloat while keeping intrinsic bloat low as it slows down non-straggler pipelines. Therefore, we report (1) intrinsic bloat reduction without stragglers and (2) intrinsic + extrinsic bloat reduction with stragglers.
@@
  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s time–energy frontier against that of Zeus (§6.4).